# Categorical Time-Reversal Symmetries

### 30 Mar 2026 v1

@@ Appendix A More on the Semi-direct Product Categories $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$
Here we will establish the Morita equivalence between an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$ and the corresponding semi-direct product Galois-real fusion category $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$, where $\mathcal{C}:=\mathcal{D}_{\mathbb{R}}\boxtimes_{\mathbb{R}}\mathsf{Vec}_{\mathbb{C}}$ is the complexification of $\mathcal{D}$. The physical meaning of this Morita equivalence is discussed in Section IV.4: if $\mathcal{C}$ is a (complex) fusion category of topological defects in a 2D theory, and the 2D theory is invariant under a ${\mathbb{Z}}_{2}^{T}$-action, then the ${\mathbb{Z}}_{2}^{T}$-invariant topological defects form a real form $\mathcal{D}_{\mathbb{R}}$ of $\mathcal{C}$, and spontaneously breaking ${\mathbb{Z}}_{2}^{T}$ results a theory with topological defects forming $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$. We formulate this as the following theorem.

### 02 Apr 2026 v2

@@ Appendix A More on the Semi-direct Product Categories $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$
Here we will establish the Morita equivalence between an $\mathbb{R}$-real fusion category $\mathcal{D}_{\mathbb{R}}$ and the corresponding semi-direct product Galois-real fusion category $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$, where $\mathcal{C}:=\mathcal{D}_{\mathbb{R}}\boxtimes_{\mathbb{R}}\mathsf{Vec}_{\mathbb{C}}$ is the complexification of $\mathcal{D}$. The physical meaning of this Morita equivalence is discussed in Section IV.4: if $\mathcal{C}$ is a (complex) fusion category of topological defects in a 2D theory, and the 2D theory is invariant under a ${\mathbb{Z}}_{2}^{T}$-action, then the ${\mathbb{Z}}_{2}^{T}$-invariant topological defects form a real form $\mathcal{D}_{\mathbb{R}}$ of $\mathcal{C}$, and spontaneously breaking ${\mathbb{Z}}_{2}^{T}$ results in a theory with topological defects forming $\mathcal{C}\rtimes{\mathbb{Z}}_{2}^{T}$. We formulate this as the following theorem.